# MobileWorld: Benchmarking Autonomous Mobile Agents in Agent-User Interactive, and MCP-Augmented Environments

### v1

@@ Abstract
To bridge this gap, we introduce MobileWorld, a substantially more challenging benchmark designed to better reflect real-world mobile usage, comprising 201 tasks across 20 applications, while maintaining the same rigorous level of reproducible evaluation as AndroidWorld. The difficulty of MobileWorld is twofold. First, it emphasizes long-horizon tasks with cross-application interactions: MobileWorld requires nearly twice as many task-completion steps on average (27.8 vs. 14.3) and includes far more multi-application tasks (62.2% vs. 9.5%) compared to AndroidWorld. Second, MobileWorld extends beyond standard GUI manipulation by introducing novel task categories, including agent–user interaction and Model Context Protocol (MCP)-augmented tasks.
To ensure robust evaluation, we provide snapshot-based container environment and precise functional verifications, including backend database inspection and task callback APIs. We further develop a planner-executor agentic framework with extended action spaces to support user interactions and MCP calls. Our results reveal a sharp performance drop compared to AndroidWorld, with the best agentic framework and end-to-end model achieving 51.7% and 20.9% success rates, respectively, highlighting ample headroom for future research. Our analysis further shows that current models struggle significantly with user interaction and MCP calls. By identifying these core research gaps, we offer a strategic roadmap toward more robust, next-generation mobile intelligence.
@@ 1 Introduction
Beyond saturation, current mobile agent benchmarks exhibit several fundamental limitations. First, existing tasks lack the complexity reflecting real-world mobile usage. Prior benchmarks typically feature short-horizon tasks confined within single applications, whereas practical mobile assistance frequently demands long-horizon planning and cross-application workflows. Second, existing benchmarks often assume fully-specified instructions, neglecting realistic scenarios where users provide vague or incomplete requests. In real-world deployments, agents must proactively engage in clarification dialogues to acquire missing information. Third, existing mobile-use benchmarks overlook the growing importance of external tool invocation. The Model Context Protocol (MCP) (anthropic_mcp_intro) has rapidly emerged as a standardized interface exposing agents to external tools beyond traditional GUI operations. The community is actively debating the integration of GUI operations alongside MCP tools (wang2025ui), recognizing that next-generation mobile agents must seamlessly combine interface manipulation with external tool invocations to solve tasks that extend beyond traditional smartphone capabilities. However, no existing benchmark evaluates this hybrid execution paradigm.
@@
Figure 2: Beyond traditional GUI-only tasks, MobileWorld includes agent–user interaction tasks and MCP-augmented tasks, each with distinct deterministic evaluation strategies. Left: An example of an agent–user interaction task, in which the agent must proactively request clarification from a simulated user when encountering incomplete information. A GPT-4.1–based simulated user agent is then triggered to provide the requested information, which is embedded in its system prompt. Task completion is verified through the application’s callback cache. Right: An example of an MCP-augmented task, where the agent is initialized with a list of GitHub MCP tools and selects the appropriate tool to retrieve README content from a GitHub repository before completing the task via GUI operations. Task completion is verified through backend database inspection.
@@
Our fourth contribution is a reproducible, end-to-end evaluation infrastructure featuring deterministic verification. As summarized in Table˜1, prior benchmarks for commercial applications — such as Gmail or YouTube — suffer from inherent limitations, including mandatory authentication protocols and opaque internal states (chai2025a3; chen2024spa; yang2025probench; yan2025stepguitechnicalreport). These restrictions frequently necessitate "MLLM-as-a-judge" evaluation, which introduces stochasticity and noise into the assessment process. We address these challenges by substituting proprietary services with production-grade, self-hosted open-source alternatives (e.g., Mattermost as a Slack alternative). By modifying source code and gaining direct access to backend databases, we establish a fully observable and controlled environment. Our framework employs a multi-faceted validation suite — comprising textual response verification, backend database inspection, local storage analysis, and application-specific callbacks — to ensure absolute determinism. Finally, by leveraging Docker-based Android Virtual Devices (AVDs) and snapshot-based state management, we provide a standardized, "push-button" evaluation protocol designed for rigorous community benchmarking.
@@ 3.4 Data Statistics
We summarize key statistics of MobileWorld in Figure˜4 and Table˜5, highlighting four core dimensions. (1) Tasks span diverse domains, including communication, productivity, and navigation. Approximately 95% of tasks involve third-party applications, aligning with authentic mobile usage patterns and ensuring the benchmark’s real-world relevance. (2) The benchmark emphasizes cross-app complexity: 62.2% of tasks require coordination between multiple apps, and 12.4% involve three or more. (3) Agent–user interaction tasks and MCP-augmented tasks account for 42.3% of the dataset. These tasks go beyond pure GUI navigation by requiring dynamic user engagement and external tool use. (4) Multiple verification methods are used to assess task success: 47.3% of tasks rely on self-hosted database verification, and 36.8% use local storage inspection, highlighting the importance of apps with fully controlled backends.
@@ 4.4 Main Results
Second, there exists a substantial performance gap between agentic frameworks and end-to-end GUI-specialized models. Agentic frameworks leveraging frontier LLMs (e.g., GPT-5, Claude-4.5-Sonnet and Gemini-3-Pro) consistently outperform end-to-end models by a large margin. The best agentic model achieves an overall success rate of 51.7%, compared to only 20.9% for the best end-to-end model (Doubao-1.5-UI-TARS). This gap stems from two key limitations of current end-to-end approaches: (1) insufficient capability in handling the complex reasoning and cross-app coordination required by MobileWorld, and (2) a lack of support for agent–user interaction and MCP tool invocation.
Third, agent-user interaction and MCP-augmented tasks prove particularly challenging and reveal critical capability gaps. On agent–user interaction tasks, Qwen3-VL and GELab-Zero achieve scores below 10%, while Doubao-1.5-UI-TARS achieves only 32.4%. This highlights the difficulties in recognizing when user clarification is needed. In contrast, GPT-5 demonstrates strong performance (62.2%) on these tasks, suggesting that advanced reasoning capabilities are essential for effective human-agent collaboration. Similarly, MCP-augmented tasks expose fundamental limitations in tool integration. Qwen3-VL is unable to properly utilize MCP tools when required to coordinate their invocation with GUI actions. On the other hand, among agentic frameworks, the best performance reaches only 51.6% (GPT-5). This reveals that current models struggle to effectively orchestrate between GUI actions and external tool invocations.
@@ 4.5 In-depth Analysis
The Average Queries and UIQ metrics jointly reveal how models manage agent–user interaction tasks. GPT-5 demonstrates the most effective clarification behavior, with an average of 1.11 queries per interaction task — corresponding to its highest user-interaction success rate (62.2%) and UIQ (0.40). Claude-4.5-Sonnet follows, averaging 0.76 queries and achieving a UIQ of 0.25. In contrast, Gemini-3-Pro got a lower UIQ (0.19) and asks very few queries on average (0.36), even though it performs well on GUI-only tasks (55.6%). This suggests it under-utilizes user clarification even when beneficial. Notably, UIQ scores are consistently lower than User-Interaction SR across all models, indicating the presence of unnecessary ask_user operations that inflate query counts without contributing to task success. In particular for Doubao-1.5-UI-TARS, a User-Interaction SR of 32.4% was achieved with 1.22 average queries, but the model obtains a lower UIQ of only 0.13. This gap suggests that Doubao-1.5-UI-TARS frequently issues redundant or ineffective clarification requests.

### v2

# MobileWorld: Benchmarking Autonomous Mobile Agents in Agent-User Interactive and MCP-Augmented Environments

@@ Abstract
To bridge these gaps, we introduce MobileWorld, a substantially more challenging benchmark designed to reflect real-world usage through 201 tasks across 20 applications. MobileWorld derives its difficulty from an emphasis on long-horizon, cross-application workflows, requiring nearly twice as many completion steps on average (27.8 vs. 14.3) and featuring a significantly higher proportion of multi-app tasks (62.2% vs. 9.5%) than AndroidWorld. To overcome the limitations of existing environments, MobileWorld achieves a balance between production-grade utility and reproducible evaluation by utilizing open-source alternatives to industry standards (e.g., Mattermost for Slack). This approach enables a fully observable and controlled environment through source code modification and direct backend database access for precise verification. Furthermore, MobileWorld extends beyond standard GUI manipulation by introducing novel task categories, including agent-user interaction and Model Context Protocol (MCP)-augmented tasks, providing a robust framework for evaluating agents in user-aware, hybrid-tool scenarios.
To facilitate evaluation, we develop a planner-executor agentic framework with extended action spaces to support user interactions and MCP calls. Our results reveal a sharp performance drop compared to AndroidWorld, with the best agentic framework and end-to-end model achieving 51.7% and 20.9% success rates, respectively, highlighting ample headroom for future research. Our analysis further shows that current models struggle significantly with user interaction and MCP calls. By identifying these core research gaps, we offer a strategic roadmap toward next-generation mobile intelligence.
@@ 1 Introduction
Beyond saturation, current mobile agent benchmarks exhibit several fundamental limitations. (1) Existing tasks lack the complexity to reflect real-world mobile usage. Prominent benchmarks like AndroidWorld omits critical application categories—such as e-commerce and enterprise communication—that are central to daily mobile workflows. Furthermore, prior benchmarks typically feature short-horizon tasks confined within single applications, whereas practical mobile assistance frequently demands long-horizon planning and cross-application workflows. (2) Existing benchmarks often assume fully-specified instructions, neglecting realistic scenarios where users provide vague or incomplete requests. In real-world deployments, agents must proactively engage in clarification dialogues to acquire missing information. (3) Existing benchmarks overlook the growing importance of external tool invocation. The Model Context Protocol (MCP) (anthropic_mcp_intro) has rapidly emerged as a standardized interface exposing agents to external tools beyond traditional GUI operations. The community is actively debating the integration of GUI operations alongside MCP tools (wang2025ui), recognizing that next-generation mobile agents must seamlessly combine interface manipulation with external tool invocations to solve tasks that extend beyond traditional smartphone capabilities. However, no existing benchmark evaluates this hybrid execution paradigm.
@@
Figure 2: Beyond traditional GUI-only tasks, MobileWorld includes agent-user interaction tasks and MCP-augmented tasks, each with distinct deterministic evaluation strategies. Left: An example of an agent-user interaction task, in which the agent must proactively request clarification from a simulated user when encountering incomplete information. A GPT-4.1-based simulated user agent is then triggered to provide the requested information, which is embedded in its system prompt. Task completion is verified through the application’s callback cache. Right: An example of an MCP-augmented task, where the agent is initialized with a list of GitHub MCP tools and selects the appropriate tool to retrieve README content from a GitHub repository before completing the task via GUI operations. Task completion is verified through backend database inspection.
@@
Our fourth contribution is an evaluation infrastructure that strikes a balance between reproducible, deterministic verification and a broad coverage of essential real-world applications. As summarized in Table˜1, prior benchmarks for commercial applications — such as Gmail or YouTube — suffer from inherent limitations, including mandatory authentication protocols and opaque internal states (chai2025a3; chen2024spa; yang2025probench; yan2025stepguitechnicalreport). These restrictions frequently necessitate "MLLM-as-a-judge" evaluation, which introduces stochasticity and noise into the assessment process. We address these challenges by substituting proprietary services with production-grade, self-hosted open-source alternatives (e.g., Mattermost as a Slack alternative). By modifying source code and gaining direct access to backend databases, we establish a fully observable and controlled environment. Our framework employs a multi-faceted validation suite — comprising textual response verification, backend database inspection, local storage analysis, and application-specific callbacks — to ensure absolute determinism. Finally, by leveraging Docker-based Android Virtual Devices (AVDs) and snapshot-based state management, we provide a standardized, "push-button" evaluation protocol designed for rigorous community benchmarking.
@@ 3.4 Data Statistics
We summarize key statistics of MobileWorld in Figure˜4 and Table˜5, highlighting four core dimensions. (1) Tasks span diverse domains, including communication, productivity, and navigation. Approximately 95% of tasks involve third-party applications, aligning with authentic mobile usage patterns and ensuring the benchmark’s real-world relevance. (2) The benchmark emphasizes cross-app complexity: 62.2% of tasks require coordination between multiple apps, and 12.4% involve three or more. (3) Agent-user interaction tasks and MCP-augmented tasks account for 42.3% of the dataset. These tasks go beyond pure GUI navigation by requiring dynamic user engagement and external tool use. (4) Multiple verification methods are used to assess task success: 47.3% of tasks rely on self-hosted database verification, and 36.8% use local storage inspection, highlighting the importance of apps with fully controlled backends.
@@ 4.4 Main Results
Second, there exists a substantial performance gap between agentic frameworks and end-to-end GUI-specialized models. Agentic frameworks leveraging frontier LLMs (e.g., GPT-5, Claude-4.5-Sonnet and Gemini-3-Pro) consistently outperform end-to-end models by a large margin. The best agentic model achieves an overall success rate of 51.7%, compared to only 20.9% for the best end-to-end model (Doubao-1.5-UI-TARS). This gap stems from two key limitations of current end-to-end approaches: (1) insufficient capability in handling the complex reasoning and cross-app coordination required by MobileWorld, and (2) a lack of support for agent-user interaction and MCP tool invocation.
Third, agent-user interaction and MCP-augmented tasks prove particularly challenging and reveal critical capability gaps. On agent-user interaction tasks, Qwen3-VL and GELab-Zero achieve scores below 10%, while Doubao-1.5-UI-TARS achieves only 32.4%. This highlights the difficulties in recognizing when user clarification is needed. In contrast, GPT-5 demonstrates strong performance (62.2%) on these tasks, suggesting that advanced reasoning capabilities are essential for effective human-agent collaboration. Similarly, MCP-augmented tasks expose fundamental limitations in tool integration. Qwen3-VL is unable to properly utilize MCP tools when required to coordinate their invocation with GUI actions. On the other hand, among agentic frameworks, the best performance reaches only 51.6% (GPT-5). This reveals that current models struggle to effectively orchestrate between GUI actions and external tool invocations.
@@ 4.5 In-depth Analysis
The Average Queries and UIQ metrics jointly reveal how models manage agent-user interaction tasks. GPT-5 demonstrates the most effective clarification behavior, with an average of 1.11 queries per interaction task — corresponding to its highest user-interaction success rate (62.2%) and UIQ (0.40). Claude-4.5-Sonnet follows, averaging 0.76 queries and achieving a UIQ of 0.25. In contrast, Gemini-3-Pro got a lower UIQ (0.19) and asks very few queries on average (0.36), even though it performs well on GUI-only tasks (55.6%). This suggests it under-utilizes user clarification even when beneficial. Notably, UIQ scores are consistently lower than User-Interaction SR across all models, indicating the presence of unnecessary ask_user operations that inflate query counts without contributing to task success. In particular for Doubao-1.5-UI-TARS, a User-Interaction SR of 32.4% was achieved with 1.22 average queries, but the model obtains a lower UIQ of only 0.13. This gap suggests that Doubao-1.5-UI-TARS frequently issues redundant or ineffective clarification requests.